# Let Me Think! A Long Chain-of-Thought Can Be Worth Exponentially Many Short Ones

### v1

@@ 4.4 Results
We have found that models trained on DFS traces exploit a long CoT budget to outperform models trained on short CoTs (those generated by Shortest-Path and Path). The models trained on short CoTs tie with the DFS trained ones on very short budgets (at relatively low accuracy), but fall behind and plateau when given a higher token budget, as if they don’t succeed early on, they don’t know how to continue (which makes sense, since they have left their training distribution). The DFS model achieves perfect evidence and decision accuracy on the tasks, while the Path and Shortest-Path models struggle with the tasks when increasing the graph’s depth; achieving 11.16% and 0.0% evidence accuracies respectively on the Bridge(5) task (See Figure 4).
@@
What scenarios at inference time lead to the failure of models trained on short CoTs? Looking more closely at the evidence accuracy of the models and their behaviour in response to an input, we find that although the Shortest-Path model is trained to take the short path from the starting node of each component to its end node, it cannot distinguish between the unexplored paths attached to the current node and gets into out-of-distribution scenarios by following the wrong paths and fails to recover from them. Therefore, given its limited CoT budget, its accuracy matches with the exponentially small probability $P(\textrm{DFS}\in D(\textrm{Shortest-Path}))=\frac{1}{3\times 4^{d-1}}$ of a DFS trace that randomly chooses an unvisited adjacent node at each step, traversing the shortest path (See Figure 4). This supports the assumption that the model’s limited expressivity limits its look-ahead ability, which motivated the Vertex Query Model (VQM) of Section 3.2. The Path model’s accuracy follows a similar trend, but it is slightly higher than the in-distribution exploration probability $P(\textrm{DFS}\in D(\textrm{Path}))=\frac{2^{d}}{3\times 4^{d-1}}$. The Path model’s more flexible CoT budget and its ability to follow edges allows it to recover from some of the out-of-distribution scenarios it gets into by backtracking from the deadend or backward paths.
@@
Since Figure 4 shows that sequential scaling of one chain of thought increases the accuracy significantly, we now ask: can we aggregate multiple short CoTs (either with best-of-$n$ or with majority voting) to achieve the same accuracy as one longer CoT? If so, how many short CoTs must we aggregate? We experiment by generating many short CoTs with temperature 1.0 and measuring accuracy for both majority and best-of-$n$ aggregation methods. For one run, decision accuracy is usually higher than evidence accuracy, because the model can both rely on its CoT to make a decision and if the CoT is not a valid proof it can randomly guess between the two target nodes. This also makes decision accuracy less robust when the model’s evidence accuracy is low (See Figure 11 for evidence of both behaviors in short CoT models). However, when parallel scaling, the best-of-$n$ method that uses CoTs scales better than taking majority over the decisions (See Figure 5). Hence, we report the best-of-$n$ accuracy for the experiments with parallel scaling. We find that the best-of-$n$ accuracies of the Shortest-Path and Path models follow the probability that at least one of their $n$ independent sampled CoTs succeed, with the success probability of each corresponding to the exponentially small probability of traversing the shortest path and a path respectively (See Figure 5). Therefore, we need to sample an exponential number of CoTs from these models to achieve an accuracy on par with a single CoT of our models trained on long CoTs (See Figure 13).
Figure 5: (left) Majority decision and (right) Best-of-$n$ accuracy for parallel scaling of models trained with CoT strategies for Bridge(3) task, compared to the accuracy predicted by each CoT independently meeting the evidence criteria with the probability of a DFS trace becoming the shortest path and a path respectively. Outputs are sampled with temperature 1.0. Error bars represent 95% binomial confidence intervals.
@@
Inspired by the observation that our short CoT models behave like the search models but with limited sequential CoT budget, we examine the evidence accuracy of each model with different sequential scales of CoT budget. We budget-force (MYS+, 25) the models by considering their CoTs of various maximum lengths, and find that at every sequential CoT budget, the DFS model achieves the highest evidence accuracy (See Figure 4). Then, to examine the best accuracy we could get with parallel scaling our models within a fixed sequential budget, we parallel scale the DFS model of different sequential scales. We find that sequential scaling up to a certain threshold is more effective than exponential parallel scaling (See Figure 1). Even from that threshold, scaling sequentially further is more efficient in terms of the number of tokens generated than parallel scaling (See Figure 6). In other words, for a fixed total token budget, sequential scaling always beats parallel scaling. This shows that sequential scaling to a long enough CoT budget given the model’s expressivity for a certain complexity task is necessary for the parallel scaling to work.
@@ 4.5 Experiments with Large Language Models
We also measured the performance of various LLMs on this task. Specifically, we consider the task on a bridge graph with short=3, long=9, deadend=0, depth=2. Mirroring our other experimental and theoretical results, LLMs only get trivial performance without a CoT, but when allowed a long CoT, can achieve very high performance.
We obtained similar trends with each of the three LLMs we tested, Qwen3-32BYLY+ (25), DeepSeek R1 Distill Qwen-32BGYZ+ (25), and S1-32BMYS+ (25) (see Figure 1 and Figure 7). Note that it takes roughly a thousand tokens of sequential scaling to get non-trivial accuracy, many of these tokens are used up by the LLM describing what its general approach to the problem will be, before actually executing a strategy.
Figure 7: A comparison of parallel and sequential scaling for s1-32B MYS+ (25) and qwen3-32B YLY+ (25). Note that the trend from Figure 1 is repeated. Sequential scaling is basically essential to get higher accuracy, and parallel scaling only becomes useful once sequential scaling has allowed for non-trivial performance.
@@ 5 Emergent sequential scaling with reinforcement learning
We observed that our models trained on short CoTs cannot look ahead to distinguish the correct and wrong paths and get into out-of-distribution scenarios. However, there are cases where the Path model recovers from the out-of-distribution scenario and succeeds in generating a verified CoT followed by the correct decision. This results in verified CoTs that are longer than any CoT in the training data (See Figure 4), with behaviors such as backtracking that are not present in the training data. How does reinforcing the model on its own verified CoTs, including these out-of-distribution CoTs, affect the model’s performance and reasoning behavior? In this section, we experiment this by using Self-Taught Reasoner (STaR) (ZWMG, 22), an expert iteration RL method, to fine-tune the model on its verified CoTs.
@@ Results
We find that the model’s accuracy on each task improves dramatically after a few iterations; evidence accuracy of the model pre-trained on Bridge(3) task jumping from 21.16% to 92.02% after 4 iterations. At the same time, the average length of the model’s valid CoTs and verified CoTs increases and the model learns to exploit increasingly longer CoTs after each RL iteration (See Figure 8). Moreover, we find that the model’s accuracy improves at every sequential CoT budget (See Figure 9). This gives insight into the observed phenomena of long CoT emergence during RL on reasoning tasks GYZ+ (25). RL can adapt the model’s training to its expressivity for the task, by reinforcing its own computations that result in solving the task. In our case, the model is not expressive enough to solve the task by following the CoT strategy it was trained on. However, after training on its successful generations during RL, its CoT scales sequentially to follow a longer but more simple strategy it is expressive enough to adopt.
6 Discussion, limitations, and future directions
@@
Additionally, while we make an effort to find the best models for graph connectivity with chain-of-thought (See Figure 4) in our experiments, we do not have a guarantee that these are indeed the best models that deploy chain of thought. In future work, this could be addressed by studying models learned with RL, with a penalty on the length of the chain of thought, to encourage more optimal use of the sequential scaling budget.
Finally, the Vertex Query Model that we propose to abstract the power of chain-of-thought in Section 3.2 is motivated by the globality barrier studied in ABL+ (24), and is empirically validated, but it does not have direct theoretical backing. An interesting future direction is to prove that bounded-depth transformers on graph connectivity tasks are indeed effectively restricted by this model.
@@ Acknowledgments
This work was initiated while EB, EE, EM, and PM were visiting the Simons Institute for the Theory of Computing. EB was supported by the Simons Institute as a Research Fellow at the Special Year on Large Language Models and Transformers, and also by NSF grant CCF-2106377. EE acknowledges a gift from AWS AI to Penn Engineering’s ASSET Center for Trustworthy AI. PM acknowledges support from the National Science Foundation (NSF) and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and #814639. SJ acknowledges funding support by the Center of Mathematical Sciences and Applications. EM was supported by the Kempner Institute for the Study of Natural and Artificial Intelligence, which was made possible in part by a gift from the Chan Zuckerberg Initiative Foundation. This work used the Delta system at the National Center for Supercomputing Applications through allocation TG-CIS220009 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296 (BDF+, 23).
@@ A.1 Preliminaries: expressivity of transformers
For any bounded-depth softmax-attention transformer $T:\Sigma^{*}\to\mathbb{R}^{|\Sigma|}$ and any polynomial $p(n)$, there is a function $\hat{T}:\Sigma^{*}\to\mathbb{R}^{|\Sigma|}$ in $\mathsf{TC}^{0}$ that approximates $T$ to $2^{-p(n)}$ additive error on inputs of length $n$ [^2] The $\mathsf{TC}^{0}$ circuit outputs in $\mathbb{R}^{|\Sigma|}$ is returned up to some number of bits of precision.
@@ A.2 Our result
formed by sampling $n$ tokens autoregressively from the transformer. $D_{T,0}$ is the empty string with probability $1$. For any $n\geq 1$, the distribution $D_{T,n}$ is the distribution of $[z_{1},\ldots,z_{n}]$ where $[z_{1},\ldots,z_{n-1}]\sim D_{T,n-1}$, and $z_{n}\sim\mathrm{softmax}(T([x;z_{1},\ldots,z_{n-1}]))$.
@@
Let $p_{1}(n)$ be a polynomial that we will fix later. Consider the circuit $T^{\prime}$ that upon input $[x;r_{1},\ldots,r_{p_{1}(n)}]$ where $x\in\Sigma^{n}$ and $r_{i}\in\{0,1\}^{p(n)}$, takes a majority vote over $\tilde{T}(x;r_{1}),\ldots,\tilde{T}(x;r_{p_{1}(n)})$. By a Chernoff bound, and a large enough polynomial $p_{1}(n)$, we have that for any $x\in\Sigma^{n}$ and $r\in\{0,1\}^{p(n)}$, we have
@@
By a union bound over all inputs $x\in|\Sigma|^{n}$, for any $n$ there is a random seed $[r_{1}^{*},\ldots,r_{p_{1}(n)}^{*}]$ such that
@@
    $\mathbb{P}_{z_{1},\ldots,z_{m(n)}\sim D_{T,C_{1}}(G,s,t)}[\mathrm{Majority}(z_{1,C_{1}},\ldots,z_{m(n),C_{1}})\neq\mathrm{ans}]<1/2+1/n\,.$
@@
Then by Lemma 1 there is a $\mathsf{TC}^{0}$ function $\hat{T}$ such that for all large enough $n$ and all size-$n$ problems $(G,s,t_{1},t_{2},\mathrm{ans})$, we have
@@
By Lemma 2, there is a $\mathsf{TC}^{0}$ function $\tilde{T}$ that approximates the autoregressively-applied $\hat{T}$, in the sense that there is a polynomial $\tilde{p}$ such that for any size-$n$ problem $(G,s,t_{1},t_{2},\mathrm{ans})$
@@
Since $\mathrm{Majority}$ is gate, the circuit $\tilde{T}(x;r_{1}),\ldots,\tilde{T}(x;r_{m(n)})$ is a $\mathsf{TC}^{0}$ function and so it can be derandomized by Lemma 3. Using this lemma, yields a $\mathsf{TC}^{0}$ function $\dot{T}$ such that for any size-$n$ problem $(G,s,t_{1},t_{2},\mathrm{ans})$,
@@ B.1 Separation in Vertex Query Model, Proof of Theorem 2
  • $O(L)$ queries sufficient: There is a VQM algorithm that executes $L-1$ queries and solves the $(s,t_{1},t_{2})$-connectivity problem with probability 1.
@@
For the positive result, consider the algorithm that queries $s$, then the neighbor of $s$, and so on, until it reaches the other end of the path. This takes at most $L-1$ queries, and reaches either $t_{1}$ or $t_{2}$, at which point the algorithm has enough information to return the correct answer.

### v2

@@ 4.4 Results
We have found that models trained on DFS traces exploit a long CoT budget to outperform models trained on short CoTs (those generated by Shortest-Path and Path). The models trained on short CoTs tie with the DFS trained ones on very short budgets (at relatively low accuracy), but fall behind and plateau when given a higher token budget, as if they don’t succeed early on, they don’t know how to continue (which makes sense, since they have left their training distribution). The DFS model achieves perfect evidence and decision accuracy on the tasks, while the Path and Shortest-Path models struggle with the tasks when increasing the graph’s depth; achieving 11.16% and 0.0% evidence accuracies respectively on the Bridge(5) task (See Figure˜4).
@@
What scenarios at inference time lead to the failure of models trained on short CoTs? Looking more closely at the evidence accuracy of the models and their behaviour in response to an input, we find that although the Shortest-Path model is trained to take the short path from the starting node of each component to its end node, it cannot distinguish between the unexplored paths attached to the current node and gets into out-of-distribution scenarios by following the wrong paths and fails to recover from them. Therefore, given its limited CoT budget, its accuracy matches with the exponentially small probability $P(\textrm{DFS}\in D(\textrm{Shortest-Path}))=\frac{1}{3\times 4^{d-1}}$ of a DFS trace that randomly chooses an unvisited adjacent node at each step, traversing the shortest path (See Figure˜4). This supports the assumption that the model’s limited expressivity limits its look-ahead ability, which motivated the Vertex Query Model (VQM) of Section 3.2. The Path model’s accuracy follows a similar trend, but it is slightly higher than the in-distribution exploration probability $P(\textrm{DFS}\in D(\textrm{Path}))=\frac{2^{d}}{3\times 4^{d-1}}$. The Path model’s more flexible CoT budget and its ability to follow edges allows it to recover from some of the out-of-distribution scenarios it gets into by backtracking from the deadend or backward paths.
@@
Since Figure 4 shows that sequential scaling of one chain of thought increases the accuracy significantly, we now ask: can we aggregate multiple short CoTs (either with best-of-$n$ or with majority voting) to achieve the same accuracy as one longer CoT? If so, how many short CoTs must we aggregate? We experiment by generating many short CoTs with temperature 1.0 and measuring accuracy for both majority and best-of-$n$ aggregation methods. For one run, decision accuracy is usually higher than evidence accuracy, because the model can both rely on its CoT to make a decision and if the CoT is not a valid proof it can randomly guess between the two target nodes. This also makes decision accuracy less robust when the model’s evidence accuracy is low (See Figure˜13 for evidence of both behaviors in short CoT models). However, when parallel scaling, the best-of-$n$ method that uses CoTs scales better than taking majority over the decisions (See Figure˜5). Hence, we report the best-of-$n$ accuracy for the experiments with parallel scaling. We find that the best-of-$n$ accuracies of the Shortest-Path and Path models follow the probability that at least one of their $n$ independent sampled CoTs succeed, with the success probability of each corresponding to the exponentially small probability of traversing the shortest path and a path respectively (See Figure˜5). Therefore, we need to sample an exponential number of CoTs from these models to achieve an accuracy on par with a single CoT of our models trained on long CoTs.
Figure 5: (left) Majority decision and (right) Best-of-$n$ accuracy for parallel scaling of models trained with CoT strategies for Bridge(3) task, compared to the accuracy predicted by each CoT independently meeting the evidence criteria with the probability of a DFS trace becoming the shortest path and a path respectively.
@@
Inspired by the observation that our short CoT models behave like the search models but with limited sequential CoT budget, we examine the evidence accuracy of each model with different sequential scales of CoT budget. We budget-force (Muennighoff et al., 2025) the models by considering their CoTs of various maximum lengths, and find that at every sequential CoT budget, the DFS model achieves the highest evidence accuracy (See Figure˜4). Then, to examine the best accuracy we could get with parallel scaling our models within a fixed sequential budget, we parallel scale the DFS model of different sequential scales. We find that sequential scaling up to a certain threshold is more effective than exponential parallel scaling (See Figure˜1). Even from that threshold, scaling sequentially further is more efficient in terms of the number of tokens generated than parallel scaling (See Figure˜8). In other words, for a fixed total token budget, sequential scaling always beats parallel scaling.
@@ 4.5 Experiments with large language models
We also measured the performance of various LLMs on the graph connectivity task, as well as on the AIME2024 (of America, 2024) benchmark.
Specifically, we consider the graph connectivity task on a bridge graph with short=3, long=9, deadend=0, depth=2. Mirroring our other experimental and theoretical results, LLMs only get trivial performance without a CoT, but when allowed a long CoT, they can achieve very high performance. We obtained similar trends with each of the three LLMs we tested, Qwen3-32BYang et al. (2025a), DeepSeek R1 Distill Qwen-32BGuo et al. (2025), and S1-32BMuennighoff et al. (2025) (See Figure˜1 and Figure˜6). Note that it takes roughly a thousand tokens of sequential scaling to get non-trivial accuracy, many of these tokens are used up by the LLM describing what its general approach to the problem will be, before actually executing a strategy.
Figure 6: A comparison of parallel and sequential scaling for s1-32B Muennighoff et al. (2025) and qwen3-32B Yang et al. (2025a). Note that the trend from Figure˜1 is repeated. Sequential scaling is basically essential to get higher accuracy, and parallel scaling only becomes useful once sequential scaling has allowed for non-trivial performance.
We also conducted experiments with the s1‑32B model (Muennighoff et al., 2025) on the AIME2024 (of America, 2024) dataset. The results show that sequential scaling can not be efficiently replaced by parallel scaling for this mathematical task, supporting the generalizability of our findings to real-world scenarios (See Figure˜12). While quantifying the exact trade-off between them for complex mathematical problems such as this is beyond the scope of our fundamental study, we observe that the results confirm our conclusion that sequential scaling is necessary.
@@ 5 Emergent sequential scaling with reinforcement learning
We observed that our models trained on short CoTs cannot look ahead to distinguish the correct and wrong paths and get into out-of-distribution scenarios. However, there are cases where the Path model recovers from the out-of-distribution scenario and succeeds in generating a verified CoT followed by the correct decision. This results in verified CoTs that are longer than any CoT in the training data (See Figure˜4), with behaviors such as backtracking that are not present in the training data. How does reinforcing the model on its own verified CoTs, including these out-of-distribution CoTs, affect the model’s performance and reasoning behavior? In this section, we explore this using Self‑Taught Reasoner (STaR) (Zelikman et al., 2022), an expert iteration RL method, to fine-tune the model on its verified CoTs.
@@ Results
We find that the model’s accuracy on each task improves dramatically after a few iterations; evidence accuracy of the model pre-trained on Bridge(3) task jumping from 21.16% to 92.02% after 4 iterations. At the same time, the average length of the model’s valid CoTs and verified CoTs increases and the model learns to exploit increasingly longer CoTs after each RL iteration (See Figure˜7). Moreover, we find that the model’s accuracy improves at every sequential CoT budget (See Figure˜9). This gives insight into the observed phenomena of long CoT emergence during RL on reasoning tasks Guo et al. (2025). RL can adapt the model’s training to its expressivity for the task, by reinforcing its own computations that result in solving the task. In our case, the model is not expressive enough to solve the task by following the CoT strategy it was trained on. However, after training on its successful generations during RL, its CoT scales sequentially to follow a longer but more simple strategy it is expressive enough to adopt.
6 Discussion
@@
Additionally, while we make an effort to find the best models for graph connectivity with chain-of-thought (See Figure˜4) in our experiments, we do not have a guarantee that these are indeed the best models that deploy chain of thought. In future work, this could be addressed by studying models learned with RL, with a penalty on the length of the chain of thought, to encourage more optimal use of the sequential scaling budget.
Finally, the Vertex Query Model that we propose to abstract the power of chain-of-thought in Section˜3.2 is motivated by the globality barrier studied in Abbe et al. (2024), and is empirically validated, but it does not have direct theoretical backing. An interesting future direction is to prove that bounded-depth transformers on graph connectivity tasks are indeed effectively restricted by this model.
@@ Acknowledgments
This work was initiated while EB, EE, EM, and PM were visiting the Simons Institute for the Theory of Computing. EB was supported by the Simons Institute as a Research Fellow at the Special Year on Large Language Models and Transformers, and also by NSF grant CCF-2106377. EE acknowledges a gift from AWS AI to Penn Engineering’s ASSET Center for Trustworthy AI. PM acknowledges support from the National Science Foundation (NSF), the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning, and the Office of Naval Research through awards DMS-2031883, #814639, and ONR-N000142412631. SJ acknowledges funding support by the Center of Mathematical Sciences and Applications. EM was supported by the Kempner Institute for the Study of Natural and Artificial Intelligence, which was made possible in part by a gift from the Chan Zuckerberg Initiative Foundation. This work used the Delta system at the National Center for Supercomputing Applications through allocation TG-CIS220009 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296 (Boerner et al., 2023).
@@ A.1 Preliminaries: expressivity of transformers
For any bounded-depth softmax-attention transformer $T:\Sigma^{*}\to\mathbb{R}^{|\Sigma|}$ and any polynomial $p(n)$, there is a function $\hat{T}:\Sigma^{*}\to\mathbb{R}^{|\Sigma|}$ in $\mathsf{TC}^{0}$ that approximates $T$ to $2^{-p(n)}$ additive error on inputs of length nn. [^1] The $\mathsf{TC}^{0}$ circuit outputs in $\mathbb{R}^{|\Sigma|}$ is returned up to some number of bits of precision.
@@ A.2 Our result
formed by sampling $n$ tokens autoregressively from the transformer. $D_{T,0}$ is the empty string with probability $1$. For any $n\geq 1$, the distribution $D_{T,n}$ is the distribution of $z_{1},\ldots,z_{n}]$ where $z_{1},\ldots,z_{n-1}]\sim D_{T,n-1}$, and $z_{n}\sim\mathrm{softmax}(T([x;z_{1},\ldots,z_{n-1}]))$.
@@
Let $p_{1}(n)$ be a polynomial that we will fix later. Consider the circuit $T^{\prime}$ that upon input $x;r_{1},\ldots,r_{p_{1}(n)}]$ where $x\in\Sigma^{n}$ and $r_{i}\in\{0,1\}^{p(n)}$, takes a majority vote over $\tilde{T}(x;r_{1}),\ldots,\tilde{T}(x;r_{p_{1}(n)})$. By a Chernoff bound, and a large enough polynomial $p_{1}(n)$, we have that for any $x\in\Sigma^{n}$ and $r\in\{0,1\}^{p(n)}$, we have
@@
By a union bound over all inputs $x\in|\Sigma|^{n}$, for any $n$ there is a random seed $r_{1}^{*},\ldots,r_{p_{1}(n)}^{*}]$ such that
@@
    $\mathbb{P}_{z_{1},\ldots,z_{m(n)}\sim D_{T,C_{1}}(G,s,t)}[\mathrm{Majority}(z_{1,C_{1}},\ldots,z_{m(n),C_{1}})=\mathrm{ans}]<1/2+1/n\,.$
@@
Then by Lemma 1 with precision $1/n$, and by triangle inequality, there is a $\mathsf{TC}^{0}$ function $\hat{T}$ such that for all large enough $n$ and all size-$n$ problems $(G,s,t_{1},t_{2},\mathrm{ans})$, we have
@@
By Lemma 2 again with precision $1/n$, and by a triangle inequality, there is a $\mathsf{TC}^{0}$ function $\tilde{T}$ that approximates the autoregressively-applied $\hat{T}$, in the sense that there is a polynomial $\tilde{p}$ such that for any size-$n$ problem $(G,s,t_{1},t_{2},\mathrm{ans})$
@@
Since $\mathrm{Majority}$ is a gate, the circuit $\mathrm{Majority}(\tilde{T}(x;r_{1}),\ldots,\tilde{T}(x;r_{m(n)}))$ is a $\mathsf{TC}^{0}$ function and so it can be derandomized by Lemma 3. Using this lemma, yields a $\mathsf{TC}^{0}$ function $\dot{T}$ such that for any size-$n$ problem $(G,s,t_{1},t_{2},\mathrm{ans})$,
@@ B.1 Separation in vertex query model, Proof of Theorem 2
  • $O(L)$ queries sufficient: There is a VQM algorithm that executes L1L-1 queries and solves the $(s,t_{1},t_{2})$-connectivity problem with probability 1.
@@
For the positive result, consider the algorithm that queries $s$, then the neighbor of $s$, and so on, until it reaches the other end of the path. This takes at most L1L-1 queries, and reaches either $t_{1}$ or $t_{2}$, at which point the algorithm has enough information to return the correct answer.